how do you work out this problem?
{2x+10y=-2
{6x+4y=20

Answer by checkley71(8403)

2X+10Y=-2
6X+4Y=20 MULTIPLYING EQUATION 1 BY -3 WE GET
-6X-30Y=6 NOW ADDING THESE TWO EQUATIONS WE GET
---------
-26Y=26 OR Y=-1 NOW SUBSTITUTING -1 FOR Y IN EITHER EQUATION WE GET
2X+10(-1)=-2 OR 2X-10=-2 OR 2X=8 OR X=8/2 OR X=4
PROOF SUBSTITUTING -1 FOR Y & 4 FOR X IN EQUATION TWO WE GET
6(4)+4(-1)=20 OR 24-4=20 OR 20=20